The area of a square is a 1/4 as large as the area of a triangle. The triangle has a 16 inch side and altitude H the same as the square. Draw the diagrams. Find the length of the side of the square.
Answer by checkley77(12844)

(16h/2)1/4=h^2
16h/8=h^2
8h^2=16h
8h^2-16h=0
8h(h-2)=0
h=2 answer.
area of the square=2*2=4
Proof.
(16*2/2)1/4=2*2
32/2*1/4=4
32/8=4
4=4